One angle of a parallelogram is 16º less than three times the measure of the angle next to it. The sums of the measures of the angles is 180º. Find the measures of the adjacent measures of the parallelogram. The smaller angle measures _____, and the larger angle measures ____.
Answer by stanbon(75887)

One angle of a parallelogram is 16º less than three times the measure of the angle next to it. The sums of the measures of the angles is 180º. Find the measures of the adjacent measures of the parallelogram.
One angle is "x"
Adjacent angle is "3x-16"
----
x + 3x-16 = 180
4x = 196
x = 49 degrees
3x-16 = 3*49-16 = 141 degrees
=====================
Cheers,
Stan H.





The smaller angle measures _____, and the larger angle measures ____.